# Assessing the star formation history of all-sky and part-sky 100 pc white dwarf samples

### v1

@@ 4 Results
All four star formation histories used from Fig. 7 of Roberts et al. (2025) provide a good fit to the observations, with some slight preferences overall for the constant star formation rate of Cukanovaite et al. (2023) and the star formation rate derived in Roberts et al. (2025) using the direct age method. Therefore the focus in the following sections is on what is revealed about the 100 pc samples themselves, and less on differences between assumed SFH forms.
@@ 4.1 Gentile Fusillo et al sample
Figure 4 shows the comparison of the observed $G_{\rm abs}$ and luminosity functions with the simulated samples of white dwarfs. For the luminosity function simulations, the best fitting SFH is that of Roberts et al. (2025) based on the 40 pc sample of white dwarfs, followed by Mor et al. (2019). The constant SFH (Cukanovaite et al., 2023) and that of Fantin et al. (2019) struggle to reproduce the observations, in particular around the rising slope and peak of the function. It should be noted that none of the assumed SFH forms fit particularly well around the peak, with all of them predicting a small upturn at just below $M_{\mathrm{bol}}=15$. The reason for this upturn is not known. By comparison, the Gentile Fusillo sample seems to have a small dip in the same location, which also cannot be explained.
@@
The absolute Gaia G magnitude distribution shows many of the same trends, namely an over-prediction of stars at the peak which is not seen in the observations. Again, both Roberts et al. (2025) and Mor et al. (2019) show the best match, while a constant SFH and that of Fantin et al. (2019) are noticeably poorer. The early-peaking SFH of Fantin in particular produces too many faint stars and too few bright stars when compared to the observed 100 pc sample, even accounting for uncertainties.
@@ 4.2 Gaia Catalogue of Nearby Stars sample
Due to the similarities between the GCNS and the catalogue of Gentile Fusillo (see Fig. 2), drawn from the same underlying Gaia DR3 catalogue, it is no surprise in Fig. 5 that the GCNS sample also has a best match to the star formation histories of Roberts et al. (2025) followed by Mor et al. (2019) and Cukanovaite et al. (2023). Once again, the GCNS does not show the upturn at the peak that the simulation suggests should appear.
@@
Figure 5: Similar to Fig. 4, but for the 100 pc GCNS white dwarf catalogue (Gaia Collaboration et al., 2021).
@@ 4.3 GSPC-WD sample
Figure 6: Similar to Fig. 4, but for the 100 pc GSPC-WD catalogue (Vincent et al., 2024).
Figure 6 shows that the GSPC-WD sample best fits the SFH of Roberts et al. (2025), then Mor et al. (2019), Cukanovaite et al. (2023), and finally Fantin et al. (2019). However the luminosity function does reveal the extent of the incompleteness of the sample, as described in Sect 2.1.3 and by the authors themselves. The downturn of the luminosity function constrains the age of the sample and in order to match the downturn of the Vincent sample, the population age would need to change from a default 10.6 Gyr to a maximum of 9.3 Gyr for the constant SFH, all the way down to 8.5 Gyr for the Fantin et al. (2019) formation history. This is just within the $3\sigma$ uncertainty of our population age (Roberts et al., 2025), and shows that the incompleteness of the GSPC-WD sample does indeed significantly bias against fainter and redder white dwarfs.
@@ 4.4 SDSS
The sample from Kilic et al. (2025) behaves similarly to the previous 100 pc samples in that the formation histories from Roberts et al. (2025) and Mor et al. (2019) fit the best. As with the GSPC-WD sample, the luminosity function does reveal the extent of the incompleteness of the sample, as described in Sect 2.2.1 and by the authors. The assumed age of the population needs to be reduced from 10.6 Gyr to anywhere between 9.3 to 7.5 Gyr, depending on the selected SFH, making it increasingly less helpful for constraining the underlying SFH of the 100 pc volume. Additionally, the bias of the sample is not necessarily just against very faint white dwarfs; the gradient of the drop off in both the luminosity function and the absolute Gaia G magnitude distribution does not match the gradient in any simulation, possibly due to the authors avoiding the spectroscopic follow-up of objects more likely to be featureless DC white dwarfs below 5000 K.
Figure 7: Similar to Fig. 4, but for the part-sky 100 pc SDSS white dwarf catalogue of (Kilic et al., 2025).
@@ 4.5 DESI
Figure 8: Similar to Fig. 4, but for the 100 pc DESI white dwarf catalogue (DESI Collaboration et al., 2025).
The part-sky DESI DR1 sample agrees best with the SFH we found in Roberts et al. (2025) from the 40 pc sample as well as Mor et al. (2019). The DESI sample does not suffer from the same apparent incompleteness as SDSS and GSPC-WD in either distribution: the downturn of the luminosity function is well fitted by all four star formation histories using a population age of 10.6 Gyr.
We theorise that this seemingly better completeness is in part due to the random nature of DESI’s sampling; a location is picked at random within the footprint, with uniform distribution of tiles over the sky (Schlafly et al., 2023), and the target selected there is also random, so there is no bias to brighter or fainter objects within the white dwarf sample. The DESI white dwarf survey targets objects mostly from Gentile Fusillo et al. (2019)’s sample (Cooper et al., 2023), which is established to be largely volume complete and unbiased, as discussed in previous sections. Selection is made with photometric measurements from Gaia DR2 and astrometric measurements from Gaia EDR3 and is based on equations (1)-(7) in Gentile Fusillo et al. (2019). 99 per cent of the DESI white dwarf candidates are expected to be given a fibre during the main survey but the choice of which white dwarfs will be selected is decided right before an observation is made and therefore is not predictable (Manser et al., 2024). Overall, this gives all white dwarfs an equal chance of being observed, and as such gives the sample low bias.
@@
Without an observational sample to compare against, we aim to provide a simple prediction of what to expect from the 4MOST 100PC sub-survey by the end of the survey around 2031. Our simulations account for the part-sky coverage and the 4MOST 100 pc sub-survey magnitude limit of $G\leq 19.5$, but neglect the fibre-fed spectroscopic selection function. As with DESI, 4MOST will select white dwarfs randomly from the input catalogue of Gentile Fusillo et al. (2021). The key difference is that 4MOST relies on the more complete Gaia DR3, whereas DESI used DR2. Not all white dwarfs within the footprint are expected to be selected, primarily because the significantly smaller fibre resolving power compared to Gaia excludes targets near neighbouring sources, but also because of time-pressure from other sub-surveys. However, since the white dwarf selection operates in a largely random fashion, we expect only a first-order reduction in completeness relative to a truly volume-limited sample (by a few percent up to $\approx$30 per cent), with no current evidence for colour-, magnitude-, or distance-dependent biases.
Figure 10 shows that the sky selection makes no significant difference in the luminosity function within errors, which means that 4MOST could be used alongside all-sky samples without the need to account for biases based on the smaller sky area observed. This conclusion is bolstered by the similar results found with the DESI DR1 sample in the previous section. The magnitude limit of the 4MOST 100PC sub-survey does cause some incompleteness at the faint end of the luminosity function, as shown in the bottom panel of Fig. 10. This suggests that accurately determining the age of the Galactic disc will remain challenging in the MOS survey era. In addition to the model age systematics discussed by Cukanovaite et al. (2023); Roberts et al. (2025), the oldest and faintest white dwarfs will be missed even within the 100 pc sample, meaning that understanding the faint end of the luminosity function will be benefit more from deeper surveys, such as Vera Rubin (Fantin et al., 2020). Nevertheless, Fig. 11 demonstrates that 4MOST will be able to distinguish within error bars between the four forms of star formation used in this work.
@@
Figure 11: Similar to Fig. 4, but for the predicted 100 pc 4MOST white dwarf catalogue (de Jong et al., 2019).
@@ 5 Discussion and Conclusion
It is clear from the previous sections that the best fitting SFH does not depend on the specific Gaia 100 pc selection or sample of white dwarfs used: across all samples tested, the late peaking SFH of Roberts et al. (2025) fits best and provides consistency with the 40 pc sample of white dwarfs. This is reassuring: the smaller 40 pc sample with full spectroscopic follow-up is representative of the larger volume 100 pc samples, and potentially even larger volume samples still. It is also reassuring that part-sky samples, such as DESI and 4MOST, or samples with no medium-resolution spectroscopic coverage, such as Gaia DR3 white dwarf catalogues or future Vera Rubin catalogues, can constrain the local SFH with high accuracy. Previous results for the SFH and age of the population using part-sky samples or samples lacking spectroscopic information or both such as Rowell (2013) or Fantin et al. (2019) therefore also should not be biased by these factors and can be compared more easily. Finally, we note that a key limitation of MOS surveys is their magnitude limit, which excludes the coolest white dwarfs even within 100 pc and therefore prevents significant improvement in constraining the faint end of the luminosity function, or age of the Galactic disc, beyond what is already achievable within $\approx$40 pc.
The results also show that the assumed scale height of the Galaxy and age-velocity dispersion relation, which become relatively more important as we get further away from the Galactic disc, have been modelled accurately enough for consistency within 100 pc. As described in Roberts et al. (2025), the age dependent scale height of the Galaxy is modelled as $h=10.71(t_{\rm total})+65$ where $t_{\rm total})$ is the total age of a star. The age dependent probability of having left the sample is then $p({\rm left})=1-65/h$. However, this would need to be confirmed for larger volumes.
We show that an early peaking star formation rate such as Roberts et al. (2025) or Mor et al. (2019), or a constant star formation rate, are the best fits overall to the 100 pc samples investigated here. While the SFH derived in Roberts et al. (2025) is the best fitting SFH, the large systematics (detailed in Table 2) mean that it is statistically impossible to say which of those three should be favoured. Therefore, we advise that a constant star formation rate can be used as a simple but effective assumption (Cukanovaite et al., 2023). We also remind the reader that due to the age-velocity dispersion relation, a constant star formation does not result in a uniform stellar age distribution at present day in a volume-complete sample. Furthermore, the age distribution of white dwarfs is further biased by the missing main-sequence stars (Roberts et al., 2025).
@@
This work only focusses on white dwarfs within the 100 pc volume and only single white dwarfs within a certain mass range (primarily to exclude binary products and unconfirmed unresolved double degenerate white dwarfs). There is much scope, especially with the 4MOST 100PC sample, to both explore other stellar populations such as the much more numerous main sequence stars, which have also previously been used to investigate the SFH (Mor et al., 2019) and binary white dwarfs. The 100PC 4MOST sample could reveal information on binary mass transfer, mass ratios, ages and many other considerations so far not included in the simulations used here and in Roberts et al. (2025). Improving our understanding of these inputs will provide more accurate SFH and age determinations in the future.
@@ Acknowledgements
The observational data used in this article are published in McCleery et al. (2020); Gentile Fusillo et al. (2021); DESI Collaboration et al. (2025); O’Brien et al. (2024) and Kilic et al. (2025). The results derived in this article will be shared on a reasonable request to the corresponding author.
Appendix A: Direct age calculations
Direct Age Calculations
@@
where $t_{\mathrm{total}}$ is the total age of the white dwarf or the time it formed before the present day, $t_{\mathrm{WD}}$ is the white dwarf cooling time and includes any delays due to crystallisation, $t_{\mathrm{MS+GP}}$ is the main sequence and giant phases lifetime of the star, and $t_{\Delta}$ is the length of any cooling delays due to the presence of a merger in the star’s history (see Roberts et al. 2025). If the inclusion of a merger delay increased the total age of the white dwarf to greater than 10.6 Gyr, the merger delay was ignored in the calculation of the total age.
@@
As in the technique of the luminosity function, calculating the cooling age requires the effective temperature and radius (or equivalently mass through the mass-radius relation) of the white dwarf, while calculating the main-sequence lifetime requires the white dwarf mass (then processed through the initial-final mass relation). This requires knowing the atmospheric composition of the white dwarf. Assumptions must be made for the Gentile Fusillo et al. (2021) and GCNS (Gaia Collaboration et al., 2021) samples. We assume a hydrogen composition for stars without composition information (see Table 1).
@@
In Fig. 12, the directly calculated SFH of the Gentile Fusillo et al 100 pc sample can be seen compared to that of the 40 pc sample, and the other all-sky 100 pc samples considered in this work. All of these derived star formation histories have been corrected for the missing main-sequence stars, and for stars that have formed within the volume under investigation but are not inside it within the present day due to the age-velocity dispersion relation (see (Roberts et al., 2025) for details). The Gentile Fusillo et al direct age results show very good agreement with those of the 40 pc sample: a very low rate of star formation at the beginning of the Galaxy’s history, increasing to a higher rate of star formation within the last $\approx$5 Gyr. The age of the population, marked by the oldest stars in the sample, sits in the 12-13 Gyr bin but Gaia data uncertainties, combined with the large sensitivity of total age to white dwarf mass, mean that this limit has large error bars, above 2 Gyr either way.
@@
The direct age result for the GCNS sample shown in Fig. 12 follows the line of the Gentile Fusillo et al sample almost exactly. The GSPC-WD samples does show some statistically significant differences compared to the other two samples: the peak of star formation occurs slightly earlier in the Galaxy’s history, and the star formation rate appears to increase much more steeply in the 5-10 Gyr range. This is likely also due to the colour (and therefore age) dependent biases discussed in the main text. With fewer faint white dwarfs in the sample, the normalised distribution of age is marginally skewed compared to the other, less biased samples. The SDSS sample in Fig. 13 clearly shows a similar behaviour, while the DESI sample is close to the all-sky Gentile Fusillo et al result.
@@
Figure 12: Star formation history derived from the direct white dwarf age method for the all-sky 100 pc samples of Gentile Fusillo et al. (2021); Gaia Collaboration et al. (2021); Vincent et al. (2024) and the 40 pc sample (O’Brien et al., 2024; Roberts et al., 2025). Vertical errors are Poisson errors and the average spread of bin values across the 100 runs added in quadrature when systematic uncertainties and Gaia observational errors are applied
Figure 13: Similar to Fig. 12, but for the SDSS (Kilic et al., 2025) and DESI (DESI Collaboration et al., 2025) samples compared to 40 pc.

### v2

@@ 4 Results
All four star formation histories used from Fig. 7 of Roberts_2025 provide a good fit to the observations, with some slight preferences overall for the constant star formation rate of Cukanovaite_2023 and the star formation rate derived in Roberts_2025 using the direct age method. Therefore the focus in the following sections is on what is revealed about the 100 pc samples themselves, and less on differences between assumed SFH forms.
@@ 4.1 Gentile Fusillo et al sample
Figure 4 shows the comparison of the observed $G_{\rm abs}$ and luminosity functions with the simulated samples of white dwarfs. For the luminosity function simulations, the best fitting SFH is that of Roberts_2025 based on the 40 pc sample of white dwarfs, followed by Mor_2019. The constant SFH (Cukanovaite_2023) and that of Fantin_2019 struggle to reproduce the observations, in particular around the rising slope and peak of the function. It should be noted that none of the assumed SFH forms fit particularly well around the peak, with all of them predicting a small upturn at just below $M_{\mathrm{bol}}=15$. The reason for this upturn is not known. By comparison, the Gentile Fusillo sample seems to have a small dip in the same location, which also cannot be explained.
@@
The absolute Gaia G magnitude distribution shows many of the same trends, namely an over-prediction of stars at the peak which is not seen in the observations. Again, both Roberts_2025 and Mor_2019 show the best match, while a constant SFH and that of Fantin_2019 are noticeably poorer. The early-peaking SFH of Fantin in particular produces too many faint stars and too few bright stars when compared to the observed 100 pc sample, even accounting for uncertainties.
@@ 4.2 Gaia Catalogue of Nearby Stars sample
Due to the similarities between the GCNS and the catalogue of Gentile Fusillo (see Fig. 2), drawn from the same underlying Gaia DR3 catalogue, it is no surprise in Fig. 5 that the GCNS sample also has a best match to the star formation histories of Roberts_2025 followed by Mor_2019 and Cukanovaite_2023. Once again, the GCNS does not show the upturn at the peak that the simulation suggests should appear.
@@
Figure 5: Similar to Fig. 4, but for the 100 pc GCNS white dwarf catalogue (GCNS_2021).
@@ 4.3 GSPC-WD sample
Figure 6: Similar to Fig. 4, but for the 100 pc GSPC-WD catalogue (Vincent_2024).
Figure 6 shows that the GSPC-WD sample best fits the SFH of Roberts_2025, then Mor_2019, Cukanovaite_2023, and finally Fantin_2019. However the luminosity function does reveal the extent of the incompleteness of the sample, as described in Sect 2.1.3 and by the authors themselves. The downturn of the luminosity function constrains the age of the sample and in order to match the downturn of the Vincent sample, the population age would need to change from a default 10.6 Gyr to a maximum of 9.3 Gyr for the constant SFH, all the way down to 8.5 Gyr for the Fantin_2019 formation history. This is just within the $3\sigma$ uncertainty of our population age (Roberts_2025), and shows that the incompleteness of the GSPC-WD sample does indeed significantly bias against fainter and redder white dwarfs.
@@ 4.4 SDSS
The sample from Kilic_2025_100pc behaves similarly to the previous 100 pc samples in that the formation histories from Roberts_2025 and Mor_2019 fit the best. As with the GSPC-WD sample, the luminosity function does reveal the extent of the incompleteness of the sample, as described in Sect 2.2.1 and by the authors. The assumed age of the population needs to be reduced from 10.6 Gyr to anywhere between 9.3 to 7.5 Gyr, depending on the selected SFH, making it increasingly less helpful for constraining the underlying SFH of the 100 pc volume. Additionally, the bias of the sample is not necessarily just against very faint white dwarfs; the gradient of the drop off in both the luminosity function and the absolute Gaia G magnitude distribution does not match the gradient in any simulation, possibly due to the authors avoiding the spectroscopic follow-up of objects more likely to be featureless DC white dwarfs below 5000 K.
Figure 7: Similar to Fig. 4, but for the part-sky 100 pc SDSS white dwarf catalogue of (Kilic_2025_100pc).
@@ 4.5 DESI
Figure 8: Similar to Fig. 4, but for the 100 pc DESI white dwarf catalogue (DESI_2025).
The part-sky DESI DR1 sample agrees best with the SFH we found in Roberts_2025 from the 40 pc sample as well as Mor_2019. The DESI sample does not suffer from the same apparent incompleteness as SDSS and GSPC-WD in either distribution: the downturn of the luminosity function is well fitted by all four star formation histories using a population age of 10.6 Gyr.
We theorise that this seemingly better completeness is in part due to the random nature of DESI’s sampling; a location is picked at random within the footprint, with uniform distribution of tiles over the sky (Schlafly_2023), and the target selected there is also random, so there is no bias to brighter or fainter objects within the white dwarf sample. The DESI white dwarf survey targets objects mostly from GentileFusillo_2019’s sample (Cooper_2023), which is established to be largely volume complete and unbiased, as discussed in previous sections. Selection is made with photometric measurements from Gaia DR2 and astrometric measurements from Gaia EDR3 and is based on equations (1)-(7) in GentileFusillo_2019. 99 per cent of the DESI white dwarf candidates are expected to be given a fibre during the main survey but the choice of which white dwarfs will be selected is decided right before an observation is made and therefore is not predictable (Manser_2024). Overall, this gives all white dwarfs an equal chance of being observed, and as such gives the sample low bias.
@@
Without an observational sample to compare against, we aim to provide a simple prediction of what to expect from the 4MOST 100PC sub-survey by the end of the survey around 2031. Our simulations account for the part-sky coverage and the 4MOST 100 pc sub-survey magnitude limit of $G\leq 19.5$, but neglect the fibre-fed spectroscopic selection function. As with DESI, 4MOST will select white dwarfs randomly from the input catalogue of GentileFusillo_2021. The key difference is that 4MOST relies on the more complete Gaia DR3, whereas DESI used DR2. Not all white dwarfs within the footprint are expected to be selected, primarily because the significantly smaller fibre resolving power compared to Gaia excludes targets near neighbouring sources, but also because of time-pressure from other sub-surveys. However, since the white dwarf selection operates in a largely random fashion, we expect only a first-order reduction in completeness relative to a truly volume-limited sample (by a few percent up to $\approx$30 per cent), with no current evidence for colour-, magnitude-, or distance-dependent biases.
Figure 10 shows that the sky selection makes no significant difference in the luminosity function within errors, which means that 4MOST could be used alongside all-sky samples without the need to account for biases based on the smaller sky area observed. This conclusion is bolstered by the similar results found with the DESI DR1 sample in the previous section. The magnitude limit of the 4MOST 100PC sub-survey does cause some incompleteness at the faint end of the luminosity function, as shown in the bottom panel of Fig. 10. This suggests that accurately determining the age of the Galactic disc will remain challenging in the MOS survey era. In addition to the model age systematics discussed by Cukanovaite_2023; Roberts_2025, the oldest and faintest white dwarfs will be missed even within the 100 pc sample, meaning that understanding the faint end of the luminosity function will be benefit more from deeper surveys, such as Vera Rubin (Fantin_2020). Nevertheless, Fig. 11 demonstrates that 4MOST will be able to distinguish within error bars between the four forms of star formation used in this work.
@@
Figure 11: Similar to Fig. 4, but for the predicted 100 pc 4MOST white dwarf catalogue (4MOST).
@@ 5 Discussion and Conclusion
It is clear from the previous sections that the best fitting SFH does not depend on the specific Gaia 100 pc selection or sample of white dwarfs used: across all samples tested, the late peaking SFH of Roberts_2025 fits best and provides consistency with the 40 pc sample of white dwarfs. This is reassuring: the smaller 40 pc sample with full spectroscopic follow-up is representative of the larger volume 100 pc samples, and potentially even larger volume samples still (although larger volumes will need to factor in stellar populations like halo white dwarfs). It is also reassuring that part-sky samples, such as DESI and 4MOST, or samples with no medium-resolution spectroscopic coverage, such as Gaia DR3 white dwarf catalogues or future Vera Rubin catalogues, can constrain the local SFH with high accuracy. Previous results for the SFH and age of the population using part-sky samples or samples lacking spectroscopic information or both such as Rowell_2013 or Fantin_2019 therefore also should not be biased by these factors and can be compared more easily. Finally, we note that a key limitation of MOS surveys is their magnitude limit, which excludes the coolest white dwarfs even within 100 pc and therefore prevents significant improvement in constraining the faint end of the luminosity function, or age of the Galactic disc, beyond what is already achievable within $\approx$40 pc.
The results also show that the assumed scale height of the Galaxy and age-velocity dispersion relation, which become relatively more important as we get further away from the Galactic disc, have been modelled accurately enough for consistency within 100 pc. As described in Roberts_2025, the age dependent scale height of the Galaxy is modelled as $h=10.71(t_{\rm total})+65$ where $t_{\rm total}$ is the total age of a star. The age dependent probability of having left the sample is then $p({\rm left})=1-65/h$. However, this would need to be confirmed for larger volumes.
We show that a late peaking star formation rate such as Roberts_2025 or Mor_2019, or a constant star formation rate, are the best fits overall to the 100 pc samples investigated here. While the SFH derived in Roberts_2025 is the best fitting SFH, the large systematics (detailed in Table 2) mean that it is statistically impossible to say which of those three should be favoured. Therefore, we advise that a constant star formation rate can be used as a simple but effective assumption (Cukanovaite_2023). We also remind the reader that due to the age-velocity dispersion relation, a constant star formation does not result in a uniform stellar age distribution at present day in a volume-complete sample. Furthermore, the age distribution of white dwarfs is further biased by the missing main-sequence stars (Roberts_2025).
@@
This work only focusses on white dwarfs within the 100 pc volume and only single white dwarfs within a certain mass range (primarily to exclude binary products and unconfirmed unresolved double degenerate white dwarfs). There is much scope, especially with the 4MOST 100PC sample, to both explore other stellar populations such as the much more numerous main sequence stars, which have also previously been used to investigate the SFH (Mor_2019) and binary white dwarfs. The 100PC 4MOST sample could reveal information on binary mass transfer, mass ratios, ages and many other considerations so far not included in the simulations used here and in Roberts_2025. Improving our understanding of these inputs will provide more accurate SFH and age determinations in the future.
@@ Acknowledgements
The observational data used in this article are published in McCleery_2020; GentileFusillo_2021; DESI_2025; OBrien_2024 and Kilic_2025_100pc. The results derived in this article will be shared on a reasonable request to the corresponding author.
Appendix A Direct age calculations
@@
where $t_{\mathrm{total}}$ is the total age of the white dwarf or the time it formed before the present day, $t_{\mathrm{WD}}$ is the white dwarf cooling time and includes any delays due to crystallisation, $t_{\mathrm{MS+GP}}$ is the main sequence and giant phases lifetime of the star, and $t_{\Delta}$ is the length of any evolutionary delays due to the presence of a merger in the star’s history (see Roberts_2025). If the inclusion of a merger delay increased the total age of the white dwarf to greater than 10.6 Gyr, the merger delay was ignored in the calculation of the total age.
@@
As in the technique of the luminosity function, calculating the cooling age requires the effective temperature and radius (or equivalently mass through the mass-radius relation) of the white dwarf, while calculating the main-sequence lifetime requires the white dwarf mass (then processed through the initial-final mass relation). This requires knowing the atmospheric composition of the white dwarf. Assumptions must be made for the GentileFusillo_2021 and GCNS (GCNS_2021) samples. We assume a hydrogen composition for stars without composition information (see Table 1).
@@
In Fig. 12, the directly calculated SFH of the Gentile Fusillo et al 100 pc sample can be seen compared to that of the 40 pc sample, and the other all-sky 100 pc samples considered in this work. All of these derived star formation histories have been corrected for the missing main-sequence stars, and for stars that have formed within the volume under investigation but are not inside it within the present day due to the age-velocity dispersion relation (see Roberts_2025 for details). The Gentile Fusillo et al direct age results show very good agreement with those of the 40 pc sample: a very low rate of star formation at the beginning of the Galaxy’s history, increasing to a higher rate of star formation within the last $\approx$5 Gyr. The age of the population, marked by the oldest stars in the sample, sits in the 12-13 Gyr bin but Gaia data uncertainties, combined with the large sensitivity of total age to white dwarf mass, mean that this limit has large error bars, above 2 Gyr either way.
@@
The direct age result for the GCNS sample shown in Fig. 12 follows the line of the Gentile Fusillo et al sample almost exactly. The GSPC-WD sample does show some statistically significant differences compared to the other two samples: the peak of star formation occurs slightly earlier in the Galaxy’s history, and the star formation rate appears to increase much more steeply in the 5-10 Gyr range. This is likely also due to the colour (and therefore age) dependent biases discussed in the main text. With fewer faint white dwarfs in the sample, the normalised distribution of age is marginally skewed compared to the other, less biased samples. The SDSS sample in Fig. 13 clearly shows a similar behaviour, while the DESI sample is visually between the all-sky Gentile Fusillo et al result and the part-sky SDSS result.
@@
Figure 12: Star formation history derived from the direct white dwarf age method for the all-sky 100 pc samples of GentileFusillo_2021; GCNS_2021; Vincent_2024 and the 40 pc sample (OBrien_2024; Roberts_2025). Vertical errors are Poisson errors and the average spread of bin values across the 100 runs added in quadrature when systematic uncertainties and Gaia observational errors are applied
Figure 13: Similar to Fig. 12, but for the SDSS (Kilic_2025_100pc) and DESI (DESI_2025) samples compared to 40 pc.